OpenStudy (anonymous):

A pyramid has a square base that is 160 m on each side. What is the perimeter of the base in kilometers? A.0.0256 km B. 0.00064 km C. 0.64 km D. 3.2 km

OpenStudy (anonymous):

\[\frac{4*160}{1000}=\frac{16}{25}=0.64 \text{ km} \]

OpenStudy (anonymous):

The numerator is, in effect, summing up the square base lengths, The denominator is converting meters to kilometers. Thank you for the medal.